# Hermitian adjacency matrices with at most three distinct eigenvalues

### v1

@@ Abstract
We study oriented graphs whose Hermitian adjacency matrices of the second kind have few eigenvalues. We give a complete characterization of the oriented graphs with two distinct eigenvalues, showing that there are only four such graphs. We show that there are infinitely many regular tournaments with three distinct eigenvalues. We extend our main results to Hermitian adjacency matrices defined over other roots of unity.
@@ 1 Introduction
This was not the result we expected at the outset, given that analogous problems for signed graphs admits numerous infinite families and have so far evaded full characterizations. We refer the reader to the papers [3, 7, 13, 15, 16] for various infinite families of signed graphs with two distinct eigenvalues.
In the remainder of this section we give some context for the problem and the required preliminary definitions. In Section 2 we prove Theorem 1.1. In Section 3 we show that, in counterpoint to Theorem 1.1, there are infinitely many directed graphs with exactly three distinct $H_{\omega}$-eigenvalues. In particular, we show that there are infinitely many regular tournaments with this property by generalizing results of [2] which leverage skew-symmetric Hadamard matrices. In Section 4 we discuss the extension of Section 2 to Hermitian adjacency matrices $H_{\sigma}$ for $\sigma$ some primitive $k$th root of unity. We show that the case $k=3$ is essentially identical to $k=6$, while the case $k=4$ is closely related to signed graphs with two eigenvalues. We conclude with a discussion for general $k$ and prove that for $k>8$ the directed edge is the only oriented graph with two $H_{\sigma}$-eigenvalues when $\sigma=\cos(\frac{2\pi}{k})+i\sin(\frac{2\pi}{k})$.
@@ 1.1 Context for the Problem
Hermitian adjacency matrices for digraphs were first studied by Guo and Mohar [5] and independently by Liu and Li [11]. In both of these papers, the authors used the primitive fourth root of unity $i$.
Subsequently, Mohar [12] introduced the variant with $\omega=\frac{1+i\sqrt{3}}{2}$ which we are using at present. More precisely, he introduced a more general definition for digraphs which specializes to our definition in the case of oriented graphs. Subsequently, Li and Yu characterized the mixed graphs with $H_{\omega}$-eigenvalues in certain small intervals, [10]. Together with Zhou [18], they then characterized the mixed graphs with the least $H_{\omega}$-eigenvalue greater than $-\frac{3}{2}$.
All of these results belong to the broader study of spectral properties of the complex unit gain graphs defined by Reff [14]. Gain graphs with two distinct eigenvalues are related to many interesting and exceptional combinatorial objects such as distance-regular covering graphs and systems of equiangular or biangular lines in complex space, see the papers [4] and [17] for various perspectives on these connections.
@@ 1.2 Why Sixth Roots of Unity?
The reader may reasonably wonder why we, and others, have chosen to work specifically with $\omega=\frac{1+i\sqrt{3}}{2}$. First, note that there are only two primitive sixth roots of unity, $\omega$ and $\overline{\omega}$ and that $H_{\omega}(D)=H^{T}_{\overline{\omega}}(D).$ So the two choices of primitive root sixth root give cospectral matrices and we lose nothing by working specifically with $\omega$. But why sixth roots?
The motivation for this choice is presented in [12], where it is pointed out that $\omega+\overline{\omega}=1$ perfectly models the convention for mixed graphs that having two directed arcs in opposite directions is equivalent to an undirected edge. This aspect will not appear for us, since we are working with oriented graphs, but still, the fact that $\omega+\overline{\omega}$ is an integer is quite useful.
For instance, using the determinant expansion of the characteristic polynomial of $H_{\omega}(D)$ one can verify that the coefficients of this characteristic polynomial are integers for any oriented graph $D$. See Theorem 2.6 from [10] for an explicit description of these coefficients. On the other hand, let $D$ be a triangle, directed so that some vertex has in-degree 2, and let $\sigma=\cos(\frac{2\pi}{k})+i\sin(\frac{2\pi}{k})$ be a primitive $k$th root of unity. The constant term of the characteristic polynomial of $H_{\sigma}(D)$ is $\sigma+\overline{\sigma}=2\mbox{Re}(\sigma)=2\cos(\frac{2\pi}{k})$ which is an integer only if $k\in\{1,2,3,4,6\}$.
@@ 1.3 Preliminary Notation and Definitions
An oriented graph $D=(V,A)$ consists of a vertex set $V$ and a set $A$ of ordered pairs of vertices called the arcs. For each vertex $v\in V$ we require $(v,v)\notin A$ so that $D$ is loopless, and for each pair of vertices $u$ and $v$ at most one of $(u,v)$ or $(v,u)$ is an arc of $A$. We will write $u\rightarrow v$ or $v\leftarrow u$ to denote $(u,v)\in A$. Moreover, we extend the previous notation so that $x\rightarrow y\leftarrow v$ denotes “both $(x,y)\in A$ and $(v,y)\in A$”. The underlying graph of $D$ is the simple graph $G=(V,E)$ with $E=\{\{u,v\}:(u,v)\in A\mbox{ or }(v,u)\in A\}$. We will occasionally abuse notation and write $|E(D)|$ to mean $|A(D)|$ to avoid confusion with the standard notation of adjacency matrices. In an undirected graph we let $N(v)$ denote the neighbourhood of $v$ and we let $d(v)=|N(v)|$. We say that two vertices of $D$ are adjacent if they are adjacent in the underlying graph of $D$. We say that an oriented graph $D$ is connected if the underlying graph of $D$ is connected. The in-degree (respectively, out-degree) of a vertex $v$ is the size of the set $\{u:(u,v)\in A\}$ (respectively, $\{u:(v,u)\in A\}$) and is denoted $d^{-}(v)$ (respectively $d^{+}(v)$). An oriented graph is regular if all in-degrees and all out-degrees of all vertices are equal. A walk in a graph is a sequence of consecutively adjacent vertices. A $k$-walk in an oriented graph is a sequence $v_{1},v_{2},\dots,v_{k+1}$ of vertices in which the vertices $v_{i}$ and $v_{i+1}$ in each successive pair are adjacent. Note that we do not stipulate the direction of the arc connecting $v_{i}$ to $v_{i+1}$. If we do have the additional feature that each of $(v_{i},v_{i+1})\in A$, then we have a directed walk. The value of a walk $u_{0}u_{1}\cdots u_{r}$ in $D$ is $\prod_{i=1}^{r}H_{u_{i-1}u_{i}}$. A directed cycle is a regular oriented cycle.
We recall the interlacing theorem for Hermitian matrices (see, e.g. [6, p. 246]). We will use it several times.
@@ 2 Two Distinct Eigenvalues
In this section, we characterize the oriented graphs with precisely two distinct $H_{\omega}$-eigenvalues. Our main result is Theorem 1.1, where we show that there are exactly four such graphs.
For the remainder of this section, $D$ will be an oriented graph on $n$ vertices for which $H=H_{\omega}(D)$ has two distinct eigenvalues: $r$ with multiplicity $m$, and $s$ with multiplicity $n-m$. We assume $s<0<r$. Our proof of Theorem 1.1 consists of several small claims, which are organized as follows:
@@ 2.1 Irregular Oriented Graphs
Set $H=H_{\omega}(D)$. Let $d$ be the average out-degree of $D$ and note that
$d=\frac{1}{n}\sum_{v\in V(D)}d^{+}(v)=\frac{|E(D)|}{n}=\frac{1}{2n}\sum_{v\in V(D)}d(v).$
Calculating the trace of $H$ gives
$0=\operatorname{tr}(H)=mr+(n-m)s.$
So
$s=\frac{mr}{m-n}$ (1)
Now, on the one hand $\operatorname{tr}(H^{2})$ is the sum of the squares of the eigenvalues of $H$. On the other hand, $\operatorname{tr}(H^{2})$ is the sum, over all vertices of $D$, of the value of 2-walks from $v$ to $v$. There is precisely one such walk for each arc incident with $v$, and the value of each such walk is $\omega\overline{\omega}=1$, so we obtain
$2dn=\sum_{v\in V(D)}d(v)=\operatorname{tr}(H^{2})=mr^{2}+(n-m)s^{2}.$ (2)
Using (1) in (2) gives
$2dn=mr^{2}+(n-m)\left(\frac{mr}{m-n}\right)^{2}=mr^{2}\left(1+\frac{m}{n-m}\right)=\frac{mnr^{2}}{n-m}$
Thus we have,
$mr=\frac{2(n-m)d}{r}.$ (3)
Now let $j$ denote the all-ones vector of length $n$. Since the entries of $H$ are $\omega$ and $\bar{\omega}$ with $\omega+\bar{\omega}=1$ and since $r$ is the largest eigenvalue of $H$ we have
$r\geq\frac{j^{\intercal}Hj}{j^{\intercal}j}=\frac{1}{n}\sum_{i>j}(H_{i,j}+H_{j,i})=\frac{|E(D)|}{n}=d,$
with equality if and only if $D$ is regular. Using Equations (1) and (3) we can bound $s$ by
$s=\frac{mr}{m-n}=\frac{2(n-m)d}{r(m-n)}=-\frac{2d}{r}\geq-2.$
The next lemma is a Hermitian variant of Lemma 2.1 from [13].
@@
Let $H$ be the adjacency matrix of $D$. The minimal polynomial of $H$ is $(x-r)(x-s)=x^{2}-(r+s)x+rs$, so
$H^{2}-(r+s)H+rsI=0.$ (4)
Again, let $n=|V(D)|$. If $m=\frac{n}{2}$, then since $\operatorname{tr}(H)=0$ we have $r=-s$ and
@@
Now assume $m\neq\frac{n}{2}$ and pick vertices $u,v$ so that $u\rightarrow v$. We consider the $(u,v)$ entry of the matrix equation (4). The $(u,v)$ entry of $H^{2}$ is the sum of the value of all 2-walks from $u$ to $v$, hence it is equal to $a+b\omega^{2}+c\omega^{4}$ for some non-negative integers $a,b,c$.
@@
Since $01+\omega^{2}+\omega^{4}=$ 0 we may rewrite the above as
@@
Since $r$ and $s$ are the eigenvalues of a Hermitian matrix they are real, so the right-hand side is real. The left-hand side is an integer multiple of $\omega$, so we must have $b-a=0$, hence $r+s=a-c$ is an integer.
@@
If $G$ is the underlying graph of the oriented graph $D$ with exactly two distinct $H_{\omega}$-eigenvalues $r,s$, where $r=-s$, then for any pair of vertices $u,v$ in $G$, the quantity $|N(u)\cap N(v)|$ is a multiple of $3$.
@@
From the minimal polynomial of $H$ we find that $H^{2}=s^{2}I$. Hence for each pair $\{u,v\}$ of distinct vertices of $D$, the sum of the values on 2-walks from $u$ to $v$ is 0. For a pair of vertices at distance 2, the 2-walks between them take values in $\{1,\omega^{2},\omega^{4}\}$ (with two different orientations that allow for the value 1). Hence there are non-negative integers $a,b,c$ so that $a+b\omega^{2}+c\omega^{4}=0$. Since $01+\omega^{2}+\omega^{4}=$ 0, we may rewrite the above as
@@
If there is no pair of vertices at distance 2, then $G$ is complete, but since $G$ is 3-regular it must be $K_{4}$, and any pair of vertices have only two common neighbors, a contradiction. So there is a pair of vertices $\{u,v\}$ at distance 2. They have at least one common neighbour, hence they have exactly three common neighbours $\{x,y,z\}$. Now, $x$ and $y$ have common neighbours $u$ and $v$, so they must share another common neighbour. This third common neighbour cannot be $z$, because then $z$ would have degree at least 4. So there is a sixth vertex $w$ incident with both $x$ and $y$. By the same argument, $x$ and $z$ must have a third common neighbour, and since $x$ already has degree $3$, this neighbour must be $w$. Hence $G$ is a $3$-regular graph containing $K_{3,3}$ as a subgraph. Since $G$ is connected we are done. ∎
@@
Let $D$ be a connected oriented graph with two distinct $H_{\omega}$-eigenvalues $r$ and $s$, where $r>s$. If $s>-2$, then $D$ is either a single directed edge or the orientation of $K_{3,3}$ from Figure 1.
@@
Since $s>-2$, Lemma 2.2 implies $s\in\{-1,-\sqrt{2},-\sqrt{3}\}.$
If $s=-1$, then $D$ has at least two vertices. Indeed, the connected oriented graph on two vertices has two distinct eigenvalues $1$ and $-1$. However, if $D$ has more than two vertices then it contains an induced subgraph on three vertices. All connected oriented graphs on three vertices have minimum $H_{\omega}$-eigenvalue less than $-1$, so interlacing implies that no graph on more than two vertices has $s=-1$.
@@
2.2 Regular Oriented Graphs

### v2

@@ Abstract
We study oriented graphs whose Hermitian adjacency matrices of the second kind have few eigenvalues. We give a complete characterization of the oriented graphs with two distinct eigenvalues, showing that there are only four such graphs. We extend this result to mixed graphs. We show that there are infinitely many regular tournaments with three distinct eigenvalues. We extend our main results to Hermitian adjacency matrices defined over other roots of unity.
@@ 1 Introduction
This was not the result we expected at the outset, given that analogous problems for signed graphs admit numerous infinite families and have so far evaded full characterizations. We refer the reader to the papers [G17, HTW19, R22, ST91, S22] for various infinite families of signed graphs with two distinct eigenvalues.
In the remainder of this section we give some context for the problem and the required preliminary definitions. In Section 2 we prove Theorem 1.1. We then extend the characterization to mixed graphs with two eigenvalues obtaining only $C_{4}$ and trivial mixed orientations of $K_{n}$ as additional examples. In Section 3 we show that, in counterpoint to Theorem 1.1, there are infinitely many directed graphs with exactly three distinct $H_{\omega}$-eigenvalues. In particular, we show that there are infinitely many regular tournaments with this property by generalizing results of [dCGKPM92] which leverage skew-symmetric Hadamard matrices. In Section 4 we discuss the extension of Section 2 to Hermitian adjacency matrices $H_{\sigma}$ for $\sigma$ some primitive $k$th root of unity. We show that the case $k=3$ is essentially identical to $k=6$, while the case $k=4$ is closely related to signed graphs with two eigenvalues. We conclude with a discussion for general $k$ and prove that for $k>8$ the directed edge is the only oriented graph with two $H_{\sigma}$-eigenvalues when $\sigma=\cos(\frac{2\pi}{k})+i\sin(\frac{2\pi}{k})$.
@@ 1.1 Context for the Problem
Hermitian adjacency matrices for digraphs were first studied by Guo and Mohar [GM15] and independently by Liu and Li [LL15]. In both of these papers, the authors used the primitive fourth root of unity $i$.
Subsequently, Mohar [M20] introduced the variant with $\omega=\frac{1+i\sqrt{3}}{2}$ which we are using at present. More precisely, he introduced a more general definition for digraphs which specializes to our definition in the case of oriented graphs. Subsequently, Li and Yu characterized the mixed graphs with $H_{\omega}$-eigenvalues in certain small intervals, [LY21]. Together with Zhou [ZLY23], they then characterized the mixed graphs with the least $H_{\omega}$-eigenvalue greater than $\frac{3}{2}$.
All of these results belong to the broader study of spectral properties of the complex unit gain graphs defined by Reff [R11]. Gain graphs with two distinct eigenvalues are related to many interesting and exceptional combinatorial objects such as distance-regular covering graphs and systems of equiangular or biangular lines in complex space, see the papers [GLS21] and [vDW22b] for various perspectives on these connections.
@@ 1.2 Why Sixth Roots of Unity?
The reader may reasonably wonder why we, and others, have chosen to work specifically with $\omega=\frac{1+i\sqrt{3}}{2}$. First, note that there are only two primitive sixth roots of unity, $\omega$ and $\overline{\omega}$ and that $H_{\omega}(D)=H^{T}_{\overline{\omega}}(D).$ So the two choices of primitive sixth root give cospectral matrices and we lose nothing by working specifically with $\omega$. But why sixth roots?
The motivation for this choice is presented in [M20], where it is pointed out that $\omega+\overline{\omega}=1$ perfectly models the convention for mixed graphs that having two directed arcs in opposite directions is equivalent to an undirected edge. This choice has useful consequences.
For instance, using the determinant expansion of the characteristic polynomial of $H_{\omega}(D)$ one can verify that the coefficients of this characteristic polynomial are integers for any mixed graph $D$. See Theorem 2.6 from [LY21] for an explicit description of these coefficients. On the other hand, let $D$ be a triangle, directed so that some vertex has in-degree 2, and let $\sigma=\cos(\frac{2\pi}{k})+i\sin(\frac{2\pi}{k})$ be a primitive $k$th root of unity. The constant term of the characteristic polynomial of $H_{\sigma}(D)$ is $\sigma-\overline{\sigma}=-2\mbox{Re}(\sigma)=-2\cos(\frac{2\pi}{k})$ which is an integer only if $k\in\{1,2,3,4,6\}$.
@@ 1.3 Preliminary Notation and Definitions
An oriented graph $D=(V,A)$ consists of a vertex set $V$ and a set $A$ of ordered pairs of vertices called the arcs. For each vertex $v\in V$ we require $(v,v)\notin A$ so that $D$ is loopless, and for each pair of vertices $u$ and $v$ at most one of $(u,v)$ or $(v,u)$ is an arc of $A$. We will write $u\rightarrow v$ or $v\leftarrow u$ to denote $(u,v)\in A$. Moreover, we extend the previous notation so that $x\rightarrow y\leftarrow v$ denotes “both $(x,y)\in A$ and $(v,y)\in A$”. The underlying graph of $D$ is the simple graph $G=(V,E)$ with $E=\{\{u,v\}:(u,v)\in A\mbox{ or }(v,u)\in A\}$. We will occasionally abuse notation and write $|E(D)|$ to mean $|A(D)|$ to avoid confusion with the standard notation of adjacency matrices. In an undirected graph we let $N(v)$ denote the neighbourhood of $v$. We say that two vertices of $D$ are adjacent if they are adjacent in the underlying graph of $D$. We say that an oriented graph $D$ is connected if the underlying graph of $D$ is connected. The in-degree (respectively, out-degree) of a vertex $v$ is the size of the set $\{u:(u,v)\in A\}$ (respectively, $\{u:(v,u)\in A\}$) and is denoted $d^{-}(v)$ (respectively $d^{+}(v)$). The degree of $v$ is $d(v)=d^{+}(v)+d^{-}(v).$ An oriented graph is regular if all in-degrees and all out-degrees of all vertices are equal. A walk in a graph is a sequence of consecutively adjacent vertices. A $k$-walk in an oriented graph is a sequence $v_{1},v_{2},\dots,v_{k+1}$ of vertices in which the vertices $v_{i}$ and $v_{i+1}$ in each successive pair are adjacent. Note that we do not stipulate the direction of the arc connecting $v_{i}$ to $v_{i+1}$. If we do have the additional feature that each of $(v_{i},v_{i+1})\in A$, then we have a directed walk. The value of a walk $u_{0}u_{1}\cdots u_{r}$ in $D$ is $\prod_{i=1}^{r}H_{u_{i-1}u_{i}}$. A directed cycle is a regular oriented cycle.
We recall the interlacing theorem for Hermitian matrices (see, e.g. [HJ12, p. 246]). We will use it several times.
@@ 2 Two Distinct Eigenvalues
In this section, we characterize the oriented graphs with precisely two distinct $H_{\omega}$-eigenvalues. Our main result is Theorem 1.1, where we show that there are exactly four such graphs. In Theorem 2.12 we extend the result to mixed graphs.
From now through Section 2.3, $D$ will be an oriented graph on $n$ vertices for which $H=H_{\omega}(D)$ has two distinct eigenvalues: $r$ with multiplicity $m$, and $s$ with multiplicity nmn-m. We assume $s<0<r$. Our proof of Theorem 1.1 consists of several small claims, which are organized as follows:
@@ 2.1 Irregular Oriented Graphs
Set $H=H_{\omega}(D)$ and $n=|V(D)|$. Let $j$ denote the all-ones vector of length $n$. Since the entries of $H$ are $\omega$ and $\bar{\omega}$ with $\omega+\bar{\omega}=1$ and since $r$ is the largest eigenvalue of $H$ we have
$r\geq\frac{j^{\intercal}Hj}{j^{\intercal}j}=\frac{1}{n}\sum_{i>j}(H_{i,j}+H_{j,i})=\frac{|E(D)|}{n}$ (1)
with equality if and only if $D$ is regular.
Now we calculate $\operatorname{tr}(H^{2})$ in two ways. On the one hand, the minimal polynomial of $H$ is $(x-r)(x-s)=x^{2}-(r+s)x+rs$, so
$H^{2}-(r+s)H+rsI=0.$ (2)
Since $H$ has zero diagonal, this implies $\operatorname{tr}(H^{2})=-nrs.$ On the other hand, $\operatorname{tr}(H^{2})$ is the sum, over all vertices of $D$, of the value of 2-walks from $v$ to $v$. There is precisely one such walk for each arc incident with $v$, and the value of each such walk is $\omega\overline{\omega}=1$, so we obtain
$nrs=\operatorname{tr}(H^{2})=\sum_{v\in V}d(v)=2|E(D)|.$ (3)
Combining Equations (1) and (3) gives the desired bound. ∎
The next lemma is a Hermitian variant of Lemma 2.1 from [R22].
@@
Let $H=H_{\omega}(D).$ We again consider the minimal polynomial of $H$ to obtain Equation (2). This implies the underlying graph of $D$ is $|rs|$-regular.
Let $n=|V(D)|$ and let $m$ be the multiplicity of $r$. If $m=\frac{n}{2}$, then since $\operatorname{tr}(H)=0$ we have r=sr=-s and
@@
Now assume $m\neq\frac{n}{2}$ and pick vertices $u,v$ so that $u\rightarrow v$. We consider the $(u,v)$ entry of the matrix equation (2). The $(u,v)$ entry of $H^{2}$ is the sum of the values of all 2-walks from $u$ to $v$. Since $\omega^{4}=\omega^{-2}$ we may write this value as $a+b\omega^{2}+c\omega^{4}$ for some non-negative integers $a,b,c$.
@@
Since $1+\omega^{2}+\omega^{4}=0$ we may rewrite the above as
@@
Since $r$ and $s$ are the eigenvalues of a Hermitian matrix they are real, so the right-hand side is real. The left-hand side is an integer multiple of $\omega$, so we must have ba=0b-a=0, hence r+s=acr+s=a-c is an integer.
@@
If $G$ is the underlying graph of the oriented graph $D$ with exactly two distinct $H_{\omega}$-eigenvalues $r,s$, where r=sr=-s, then for any pair of vertices $u,v$ in $G$, the quantity $|N(u)\cap N(v)|$ is a multiple of $3$.
@@
From the minimal polynomial of $H$ we find that $H^{2}=s^{2}I$. Hence for each pair $\{u,v\}$ of distinct vertices of $D$, the sum of the values on 2-walks from $u$ to $v$ is 0. For a pair of distinct vertices, the 2-walks between them take values in $\{1,\omega^{2},\omega^{4}\}$ (with two different orientations that allow for the value 1). Hence there are non-negative integers $a,b,c$ so that $a+b\omega^{2}+c\omega^{4}=0$. Since $1+\omega^{2}+\omega^{4}=0$, we may rewrite the above as
@@
If there is no pair of vertices at distance 2, then $G$ is complete, but since $G$ is 3-regular it must be $K_{4}$, and any pair of distinct vertices have only two common neighbors, a contradiction. So there is a pair of vertices $\{u,v\}$ at distance 2. They have at least one common neighbour, hence they have exactly three common neighbours $\{x,y,z\}$. Now, $x$ and $y$ have common neighbours $u$ and $v$, so they must share another common neighbour. This third common neighbour cannot be $z$, because then $z$ would have degree at least 4. So there is a sixth vertex $w$ incident with both $x$ and $y$. By the same argument, $x$ and $z$ must have a third common neighbour, and since $x$ already has degree $3$, this neighbour must be $w$. Hence $G$ is a $3$-regular graph containing $K_{3,3}$ as a subgraph. Since $G$ is connected we are done. ∎
@@
Let $D$ be a connected oriented graph with two distinct $H_{\omega}$-eigenvalues $r$ and $s$, where $r>s$. If s>2s>-2, then $D$ is either a single directed edge or the orientation of $K_{3,3}$ from Figure 1.
@@
Since s>2s>-2, Lemma 2.2 implies $s\in\{-1,-\sqrt{2},-\sqrt{3}\}.$
If s=1s=-1, then $D$ has at least two vertices. Indeed, the connected oriented graph on two vertices has two distinct eigenvalues $1$ and 1-1. However, if $D$ has more than two vertices then it contains an induced subgraph on three vertices. All connected oriented graphs on three vertices have minimum $H_{\omega}$-eigenvalue less than 1-1, so interlacing implies that no graph on more than two vertices has s=1s=-1.
@@
2.2 Regular Oriented Graphs with $r\leq 2$